Description of Biorythms in own words with significance of the three curves including their relationship to one another.

Description of Biorythms as sine and cosine curves including the importance of phase and amplitude e.g. critical days

Pseudocode for the calculation of Biorythm

graphical display with significant parts of the curves labelled
Answer #1

A biorhythm is an effort to predict varied aspects of a human life through easy mathematical cycles. the speculation was developed by Wilhelm Fliess in the late nineteenth century, and was popularized within the u. s.in late Nineteen Seventies. Most scientists believe that the concept has no additional prophetical power than probability. "The theory of biorhythms could be a theory that claims our daily lives ar considerably full of rythmic cycles."



Biorhythms might not be terribly scientific, however it's true that on some days we tend to ar filled with energy whereas on different days we discover it tough simply obtaining out of bed. Biorhythms try and make a case forthis by considering the interaction between 3 cycles: emotional, physical and intellectual.

When a selected cycle is high, it means that it's straightforward to try and do well in this space (for example, if our intellectual cycle is high, we are going to do higher in examinations). however if we tend to ar within the low a part of a cycle, it's tough to try and do well (we could have additional arguments if our emotional cycle is low.)

At birth, each cycle is at 0 (neither +ve nor -ve). The cycles have different length:

  • Physical: 23 days,
  • Emotional: 28 days, and
  • Intellectual: 33 days

Interactive Biorhythms Graph

In the following interactive graph, enter your birthday at the top and you can see what is happening in your own emotional, physical and intellectual cycles. It tells you when your "good" and "not-so-good" days will occur (or have occurred in the past).

You can also use it to figure out why your friend (or mother) is crabby today!

Background

These biorythm cyclic graphs are sine curves with exactly the same amplitude, but varying period.

To obtain the graphs, we use:

  • Physical: y=sin(2πt/23​)
  • Emotional: y=sin(2πt/28​)
  • Intellectual: y=sin(2πt/33​)

The "good" and "not-so-good" days are determined by the sum of the 3 cycles (which you can see if you check the "Show composite curve" box above the graph). See: Composite Trigonometric curves for some background.

Explanation

Here's a typical composite curve for days 100 to 150:

We see that near day 130, all the cycles are against us and we will probably have a terrible day. Two weeks later, near day 145, everything is coming together nicely and we should have an excellent day.

The composite curve (in magenta color) is the total sum of the 3 cycles:

y=sin(2πt/23​)+sin(2πt/28​)+sin(2πt/33​)